### v1

@@ Results
In most quantum sensing schemes, the spin states of the NV centers are first initialized to the ||||0$\rangle$ state by a laser pulse and then transferred to a superposition state by a $\pi/2$ MW pulse. In this sensing state, alternating magnetic fields can be detected by selective manipulation of the spin state via MW pulses using pulsed (PDD) or continuous dynamical decoupling (CDD) schemes. PDD methods, such as the XY8-$N$ sequence used in this work, consists of $N$ repetitions of a block of eight $\pi$ pulses around the $x$- and $y$-axis (see Figure 1 (a)). In these conditions, a relative phase $\theta_{\mathrm{PDD}}(t_{s})=(2/\pi)\gamma\hat{B}_{\mathrm{RF}}t_{s}$ between the ||||0$\rangle$ and ||||1$\rangle$ states is accumulated only when the matching condition $\nu_{\mathrm{RF}}=1/(4\tau)$ between the RF signal frequency $\nu_{\mathrm{RF}}$ and the spacing $2\tau$ between the $\pi$ pulses is met. Here, $\gamma$ is the gyromagnetic ratio of the electron, $\hat{B}_{\mathrm{RF}}$ the RF magnetic field amplitude and $t_{s}=2N\tau$ the interrogation time during which the phase $\theta$ is accumulated. To date, this has been the method of choice in NV-based quantum sensing, as it provides optimal sensitivities for signals up to a few MHz frequency and is particularly robust to pulse imperfections due to the alternate pulse phase design degen_quantum_2017 ; barry_sensitivity_2020 ; abe_tutorial_2018 ; smits_two-dimensional_2019 ; Bucher_principles ; Liu_2022 ; bruckmaier2023imaging ; briegel_optical_2024 .
@@
In a typical CASR experiment, consecutive PDD sequences are set on resonance with the sensing RF frequency by adjusting the corresponding $\tau$ spacing according to the matching condition mentioned above. Then, the sequences are synchronized with the sensing RF signal and configured to be sensitive to the signal’s phase (see Materials degen_quantum_2017 ). This synchronization is achieved by setting the total duration of the individual PDD sequences to a multiple of the pulse spacing $\tau$. A slight deviation of the RF frequency $\nu_{\mathrm{RF}}$ with respect to the pulse sequence: $\nu_{\mathrm{PDD}}=1/(4\tau)$ results in different RF phases being recorded by each individual PDD sequence. As a consequence, the accumulated phases $\phi$, indicated by the violet and orange areas in the inset of Figure 3 (a), differ for each subsequence. Thus, the fluorescence signal (red dots in Figure 3 (a), inset) oscillates at the difference frequency $\Delta\nu=|\nu_{\mathrm{PDD}}-\nu_{\mathrm{RF}}|$ which can be Fourier transformed in order to obtain the corresponding frequency spectrum. Since the sensing time is not limited by the coherence time, the spectral resolution can be arbitrarily improved, only being limited by the clock stability of the experimental setup schmitt_submillihertz_2017 ; boss_quantum_2017 ; glenn_high-resolution_2018 .
@@
Until now, this method has been implemented exclusively using PDD protocols, resulting in the same limitation regarding the maximum detectable frequency $\nu_{\mathrm{RF}}$. Our approach exploits the phase sensitivity of the spinlock scheme. By satisfying the condition $\nu_{\mathrm{RF}}=\nu_{\mathrm{CDD}}=\Omega_{\mathrm{SL}}/(2\pi)$ and introducing a small deviation between the pulse sequence and the RF signal synchronization, a novel CDD-CASR experiment with a wider frequency detection range is enabled. In the experiments shown in Figure 3 (a), we compare the Fourier transformed signal obtained by PDD-CASR using the XY8-$N$ protocol with the CDD-CASR at different RF sensing frequencies. The detuning frequencies $\Delta\nu$ were between $150-200\,\mathrm{Hz}$. A detailed summary of experimental parameters can be found in the Methods part. As in the previous result, at a sensing frequency of $\sim 5\,\mathrm{MHz}$, the PDD-CASR exhibits a significantly higher SNR (see Figure 3 b)) than the CDD-CASR. However, it rapidly declines for increasing RF frequencies whereas the peak heights of the CDD-CASR measurements remain fairly constant over all tested frequencies, enabling sensing of RF fields up to $\sim 85\,\mathrm{MHz}$ (see Figure 3 (c)) with high frequency resolution. Although approximately eight times less sensitive than PDD at sensing frequencies as low as $\sim 5\,\mathrm{MHz}$ (see Figure 3 (b)), CDD-CASR demonstrates the ability to achieve sensitivities comparable to its counterpart at higher frequencies of around $\sim 10-15\,\mathrm{MHz}$. Importantly, it consistently performs well even at higher frequencies, proving to be a reliable method in a CASR protocol, with capabilities reaching up to 80-90 MHz.
@@ Methods
The resonator was constructed from high permittivity ceramics ($\epsilon$= 80) and was machined into the shape of a ring with an outer diameter of 4.8 mm, an inner diameter of 2 mm, and a height of 1.55 mm. It is excited by a microstrip line positioned beneath it. The desirable resonance mode for our application is TE01δ (see Figure 2 in Supplementary Note 2 of the Supplementary Information), which produces a high and homogeneous $\hat{B}_{\mathrm{RF}}$ in the center of the structure. For this mode, the resonance frequency is approximately 9.4 GHz, and the loaded quality factor $Q$ of the resonator, when critically coupled, is found to be $\sim 700$.
@@
In this experiment, the RF source generated a non-synchronized signal by keeping it unlocked in regards to the pulse sequence so that each repetition of the sequence results in a different RF phase being recorded. For each sensing frequency, the corresponding spacing between $\pi$ pulses is kept constant while the RF frequency is swept and the fluorescence signal is recorded. The phase accumulation time $t_{s}$, that is the timeframe between initial and last $\pi/2$ pulse is $5\,\upmu\mathrm{s}$ for all experiments. This, together with the dead time of $1.1\,\mathrm{ms}$ set by the laser pulse, results in a duty cycle of $\sim 0.5\%$. Such a low duty cycle is chosen to reduce heating effects of the resonator, which could negatively affect the performance of the spinlock experiment, as much as possible (see Supplementary Note 5 in the Supplementary Information).

### v2

@@ Results
In most quantum sensing schemes, the spin states of the NV centers are first initialized to the ||||0$\rangle$ state by a laser pulse and then transferred to a superposition state by a $\pi/2$ MW pulse. In this sensing state, alternating magnetic fields can be detected by selective manipulation of the spin state via MW pulses using pulsed (PDD) or continuous dynamical decoupling (CDD) schemes. PDD methods, such as the XY8-$N$ sequence used in this work, consists of $N$ repetitions of a block of eight $\pi$ pulses around the $x$- and $y$-axis (see Figure 1 (a)). In these conditions, a relative phase $\theta_{\mathrm{PDD}}(t_{s})=(2/\pi)\gamma\hat{B}_{\mathrm{RF}}t_{s}$ between the ||||0$\rangle$ and ||||1$\rangle$ states is accumulated only when the matching condition $\nu_{\mathrm{RF}}=1/(4\tau)$ between the RF signal frequency $\nu_{\mathrm{RF}}$ and the spacing $2\tau$ between the $\pi$ pulses is met. Here, $\gamma$ is the gyromagnetic ratio of the electron, $\hat{B}_{\mathrm{RF}}$ the RF magnetic field amplitude and $t_{s}=16N\tau$ the interrogation time during which the phase $\theta$ is accumulated. To date, this has been the method of choice in NV-based quantum sensing, as it provides optimal sensitivities for signals up to a few MHz frequency and is particularly robust to pulse imperfections due to the alternate pulse phase design degen_quantum_2017 ; barry_sensitivity_2020 ; abe_tutorial_2018 ; smits_two-dimensional_2019 ; Bucher_principles ; Liu_2022 ; bruckmaier2023imaging ; briegel_optical_2024 .
@@
In a typical CASR experiment, consecutive PDD sequences are set on resonance with the sensing RF frequency by adjusting the corresponding $\tau$ spacing according to the matching condition mentioned above. Then, the sequences are synchronized with the sensing RF signal and configured to be sensitive to the signal’s phase (see Materials degen_quantum_2017 ). This synchronization is achieved by setting the total duration of the individual PDD sequences to a multiple of the pulse spacing $\tau$. A slight deviation of the RF frequency $\nu_{\mathrm{RF}}$ with respect to the pulse sequence: $\nu_{\mathrm{PDD}}=1/(4\tau)$ results in different RF phases being recorded by each individual PDD sequence. As a consequence, the accumulated phases $\theta$, indicated by the violet and orange areas in the inset of Figure 3 (a), differ for each subsequence. Thus, the fluorescence signal (red dots in Figure 3 (a), inset) oscillates at the difference frequency $\Delta\nu=|\nu_{\mathrm{PDD}}-\nu_{\mathrm{RF}}|$ which can be Fourier transformed in order to obtain the corresponding frequency spectrum. Since the sensing time is not limited by the coherence time, the spectral resolution can be arbitrarily improved, only being limited by the clock stability of the experimental setup schmitt_submillihertz_2017 ; boss_quantum_2017 ; glenn_high-resolution_2018 .
@@
Until now, this method has been implemented exclusively using PDD protocols, resulting in the same limitation regarding the maximum detectable frequency $\nu_{\mathrm{RF}}$. Our approach exploits the phase sensitivity of the spinlock scheme. By satisfying the condition $\nu_{\mathrm{RF}}=\nu_{\mathrm{CDD}}=\Omega_{\mathrm{SL}}/(2\pi)$ and introducing a small deviation between the pulse sequence and the RF signal synchronization, a novel CDD-CASR experiment with a wider frequency detection range is enabled. In the experiments shown in Figure 3 (a), we compare the Fourier transformed signal obtained by PDD-CASR using the XY8-$N$ protocol with the CDD-CASR at different RF sensing frequencies. The detuning frequencies $\Delta\nu$ were between $150-200\,\mathrm{Hz}$. A detailed summary of experimental parameters can be found in the Methods part. As in the previous result, at a sensing frequency of $\sim 5\,\mathrm{MHz}$, the PDD-CASR exhibits a significantly higher SNR (see Figure 3 b)) than the CDD-CASR. However, it rapidly declines for increasing RF frequencies whereas the peak heights of the CDD-CASR measurements remain fairly constant over all tested frequencies, enabling sensing of RF fields up to $\sim 85\,\mathrm{MHz}$ (see Figure 3 (c)) with high frequency resolution. Although approximately eight times less sensitive than PDD at sensing frequencies as low as $\sim 5\,\mathrm{MHz}$ (see Figure 3 (b)), CDD-CASR demonstrates the ability to achieve sensitivities comparable to its counterpart at higher frequencies of around $\sim 10-15\,\mathrm{MHz}$. Importantly, it consistently performs well even at higher frequencies, proving to be a reliable method in a CASR protocol, with capabilities reaching up to $\sim$ 80-90 MHz.
@@ Methods
The resonator was constructed from high permittivity ceramics ($\epsilon$= 80) and was machined into the shape of a ring with an outer diameter of 4.8 mm, an inner diameter of 2 mm, and a height of 1.55 mm. It is excited by a microstrip line positioned beneath it. The desirable resonance mode for our application is TE01δ (see Figure 2 in Supplementary Note 2 of the Supplementary Information), which produces a high and homogeneous $\hat{B}_{1}$ in the center of the structure. For this mode, the resonance frequency is approximately 9.4 GHz, and the loaded quality factor $Q$ of the resonator, when critically coupled, is found to be $\sim 700$.
@@
In this experiment, the RF source generated a non-synchronized signal by keeping it unlocked in regards to the pulse sequence so that each repetition of the sequence results in a different RF phase being recorded. For each sensing frequency, the corresponding spacing between $\pi$ pulses is kept constant while the RF frequency is swept and the fluorescence signal is recorded. The phase accumulation time $t_{s}$, that is the timeframe between initial and last $\pi/2$ pulse is $5\,\upmu\mathrm{s}$ for all experiments. This, together with the dead time of $1.1\,\mathrm{ms}$ set by the laser pulse, results in a duty cycle of $\sim 0.5\%$. Such a low duty cycle is chosen to reduce heating effects of the resonator, which could negatively affect the performance of the spinlock experiment (see Supplementary Note 5 in the Supplementary Information).